The Target Copy Company is contemplating the replacement of its old printing machine with a new model costing $ 36 ,000. The old machine, which originally cost $ 33 ,000, has 3 years of expected life remaining and a current book value of $ 23 ,000 versus a current market value of $ 29 ,000. Target's corporate tax rate is 30 percent. If Target sells the old machine at market value, what is the initial after-tax outlay for the new printing machine? Since this is a cash outlay, be sure to use the - sign when writing your answer.
Answer #1

Book value of old machine = $23000

Market value = $29000

Gain on Sale = $6,000

Tax on gain = 6000*30% = $1,800

After tax salvage value of old machine = 29000-1800 = $27,200

Initial outlay of new machine = After tax salvage value of old machine - Cost of new machine

= 27200-36,000

= -$8,800